Operations Report – January 5th

Operations Report

Crew 219 Operations Report 05-01-2020

Sol: 0

Name of person filing report: Alejandro Luis Perez

Non-nominal systems: N/A

Generator : Run

Hours run : 13

From what time last night: 18:00

To what time this morning: 07:00

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before generator is run at night): 67%

Diesel Reading – 73%

Station Propane Reading- 57%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 55

Water Meter (units)- 0147363,4

Water (static Tank) (gal) – 550

Static to loft Pump used – Yes

Water in Green Hab (gal): 150 and still filling before the beginning
of the mission

Water in Science Dome (gal): 0

Toilet Tank Emptied – Yes

Diemos rover used: Still in the Workshop

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used- Yes

Hours: 122.9

Beginning Charge: 100%

End Charge: 65%

Currently Charging: Yes

Opportunity Rover used: Still in the workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

Hours: 128.3

Beginning Charge: 100%

Ending Charge: 70%

Currently Charging: Yes

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use: N/A

Oil Added? No

# hours the ATV’s were used today: 0

Notes on ATVs: N/A

Habcar used and why, where? Yes, to fill static tank and Greenhab tank
from hollow mountain Gas station

CrewCar used and why, where? Yes, to bring Crew 218 back to grand
junction. Currently in Grand junction.

General Notes and Comments: installation of the lower deck heater, but
that it won’t be operational until positioned on the right location

Summary of the internet: Not Responsive

Summary of Suit and Radios: Radio # 8 was not receiving signal. There
are currently 2 suit helmets that need repair. One has a broken head
pad panel that does not stay in place. The other is missing the
right-side connection piece for air.

Summary of Hab: The pipe connecting static tank to loft tank has to be
defrosted by using a combination of Hairdryer to the outside pipe,
warm water poured into the feed line, and re-taping the insulation
around the outside pipe.

Summary of Science Dome operations: Adjusted latching mechanism on
Science Dome door. The door now latches closed.

Summary of Ram operations: Nominal

Summary of any observatory issues: Previous Crew had replaced both the
batteries and a new Fire alarm unit. However, the alarm was
consistently going off. According to the unit the system does not
operate in conditions under 40 °F. And there is no Co alarm in there
as well.

Summary of Health and safety issues: Smoke and Co alarm have been
replaced in Greenhab. All ABC fire extinguisher are out of date, the
observatory does not have a fire extinguisher, BC fire extinguisher in
RAM has no date and is broken and three good aerosol fire
extinguishers.

Questions, concerns, and requests to Mission Support: Light bulbs
missing in 4 crew rooms, Max watts is 13.

GreenHab Report – January 5th

Crew 218 GreenHab Report 05-JAN-20

Crew GreenHab Officer: Crew 219 GreenHab Officer Cynthia Montanez

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 20.5 °C; 19.3 %

Average Max: 24.7 °C, 21.6%

Average Min: 17.3 °C, 12.6

07:56 A.M.

Floor Unit: 14 °C

Electronic: 17.4 °C

humidity 21%

Max: 23.3 °C; 22%

Min: 16.9 °C; 18%

11:50 A.M.

Floor Unit: 14 °C

Electronic: 25.1 °C

humidity 17%

Max: 25.3 °C; 22%

Min: 16.9 °C; 03%

04:48 P.M.

Floor Unit: 14 °C

Electronic: 19.2 °C

humidity 20%

Max: 25.6 °C; 21%

Min: 18.2 °C; 17%

Hours of supplemental light: Light system 7:00PM-12:00AM.

Daily water usage for crops: 0.77 Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – 67.23 Gal.

Time(s) of watering for crops:

7:56 A.M.

7:20 P.M.

SEASONAL PLANTING HISTORY:

Change to crops: None.

OTHER NOTES:

1st sprouts: N/A

Harvest: At 11:50 A.M., 2 grams of cilantro and 2 grams of marjoram
were harvested.

At 7:20 P.M., 0.25 grams of marjoram and 0.25 grams of
sage was harvested.

Other: At 7:56 A.M. there showed more flowers on the cucumber plant.
Also, at this time there were only three plants that required watering
which were icicle radish, peas, and snow peas. All other bed/
containers contained reasonable soil saturation. A removal of a dead
sweet pea plant were performed at 4:48 P.M. Additionally, a stick was
placed on one of the living sweet pea plants to make it standup.
Finally, a last watering was performed for the hanging plants, peas,
snow peas, and cucumbers, at 7:20 P.M.

Journalist Report – January 5th

Journalist Report

Robinson Raphael, Crew 219 Astronomer,

Sol 0,

The Beginning of New Potentials

Crew 219 has officially arrived on Mars! Upon arriving, we
familiarized ourselves with Crew 218 and overall had a joyous and
welcoming experience. Alongside me, the crew geologist, commander,
HSO, LSO all slept in the science dome. It was quite frightful to wake
up this morning as I was the only person in the science dome. At
first, I was so confused questioning. Did my crew leave me? Did I
oversleep?? Was I still in a dream??? After giving myself some time, I
arrived back at the Hab to see a long line to use the bathroom. Some
things never change….. I slowly found out that each crew member woke
up at different times and just left. Keith, our HSO, journeyed to the
GreenHab at 2am only to find out the heater stopped. Oh boy that story
made all of us laugh and was a wonderful way to the start the day. As
the morning progressed, we all said our goodbyes to Crew 218 and
wished them a safe journey back to Earth.

As each member of Crew 218 left, we were filled with sadness as their
company was well appreciated. At the same time, we were also filled
with motivation, as we reminisced on the research projects they were
working on. A little bit after, we all took a written quiz which both
filled our guts with anticipation and a bit of nervousness. One by one
we all handed our quizzes back to be graded and slowly waited for
results. After each quiz was graded, we let out a burst of excitement
to find out that the crew passed and was able to carry on with their
day.

The day was quite adventurous as it consisted of training with the
rovers and learning how to put on the flight suits. For the rover
training, I was paired with Alejandro and we drove all to the starting
point of Galileo Road. Man, oh man it was so cold; my face felt like a
brick. Definitely going to bring more layers next time! The scenery of
Mars was gorgeous and breathtaking that I couldn’t just help but stare
all around me. As Alejandro was driving for the first half of the
journey, we kept singing to help the time go by and it was a wonderful
moment. When it came to the flight suits, we all took pictures posing
for the camera and admiring the feel and look.

As the day began to grow darker, the whole Crew 219 spent a majority
of time doing their own thing and helping out with daily tasks such as
washing dishes, cleaning the rooms, cooking, etc. I can’t wait to wake
up to our first Sol on Mars tomorrow morning!

Sol Summary – January 5th

Crew 219 Sol Summary Report 05-01-2020

Sol:0

Summary Title:

Final Countdown to Launch

Author’s name: Dave Masaitis

Mission Status: Preparing for sim

Sol Activity Summary: Today Crew 219 spent much of the day conducting
pre-mission training, completing a proficiency quiz, inventorying food
and filling water, and of course, heating frozen pipes. Most
importantly, all maintenance attempted was completed and water got
filled, even in the loft tank. We will spend the evening polishing our
mission plan, and then we’re off to bed early, to start our simulation
fully rested.

Look Ahead Plan: Tomorrow will consist of two EVA’s to get researchers
out into areas of interest for sample collection, as well as orienting
the whole crew to our region of operations. We will also be closely
monitoring our water delivery system, though today’s work gave us some
good techniques to get water flowing despite our continuous cold
weather. We will be using down time around the campus to finish our
initial inventories. If time allows in the evening, we will also be
testing through our UAV in an attempt to make it operational after
transit.

Anomalies in work: None at present

Weather: Dominantly overcast and cold, with sporadic light flurries of snow.

Crew Physical Status: Nominal – No injuries

EVA: N/A

Reports to be file:

Sol Summary

Journalist Report

GreenHab Report

Operations Report

Photos (Including Photo of the Day)

Support Requested: See operations report